# Polarons and Exciton-Polarons in Two-Dimensional Polar Materials

### v1

@@ IV.1 Hamiltonian and unitary transformations
$V_{k}=ie\omega_{\rm t}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}\Sigma\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k}.$ (S49)
@@ IV.2 Polaron energy minimization
$|V_{k}|^{2}$ $=\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}}\frac{1}{(\varepsilon+r_{\infty}k)}.$ (S82)
@@
$\beta p$ $=\frac{\Sigma}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+2\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)\right)\frac{\hbar k\cos\varphi k{\rm d}k{\rm d}\varphi}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{3}}$
@@
$=\frac{p^{2}}{2m_{e}}(1-\beta^{2})-\frac{\Sigma}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\left(1+\frac{\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{\frac{\hbar^{2}}{m_{e}^{2}}(1-\beta)^{2}(pk\cos\varphi)^{2}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\right)k{\rm d}k{\rm d}\varphi$
@@
$\alpha_{e}=\frac{E_{0}}{\hbar\omega_{\rm t}}I_{m}(\sigma_{0},\sigma_{e,{\rm t}}),$ (S94)
@@
$\beta=\frac{\alpha_{e}}{1+\alpha_{e}},$ (S95)
@@
$E$ $=-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\alpha_{e}}.$ (S96)
We can further proceed with approximate expressions for the integral $I_{rmb}$. Applying an approximation $\sqrt{(1+\tilde{k})/(1+\sigma_{0}\tilde{k})}\approx 1/\sqrt{\sigma_{0}}$, we get
@@
$I_{\rm b}\approx\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left(\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log(\tilde{\sigma_{e,{\rm t}}})}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log(\sigma_{0})}{\sigma_{0}-1}\right),$ (S97)
where $\tilde{\sigma}_{e,{\rm t}}=\sigma_{e,{\rm t}}/\sigma_{0}^{1/4}$. Similarly,
$I_{m}\approx\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)}{8\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})^{4}}\left[2(\sigma_{0}^{6}+\tilde{\sigma}_{e,{\rm t}}^{6})-3\pi\sigma_{0}\tilde{\sigma}_{e,{\rm t}}\left(\sigma_{0}^{4}+\tilde{\sigma}_{e,{\rm t}}^{4}\right)-18\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2}-\pi)+24\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}-\tilde{\sigma}_{e,{\rm t}}^{2})\log\!\biggl{(}\sigma_{0}/\tilde{\sigma}_{e,{\rm t}}\biggr{)}\right].$ (S98)
@@ V Exciton states
$A_{\mathbf{k}}$ $=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}},\right)}$ (S133)
$B_{\mathbf{k}}$ $=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}.$ (S134)
@@
$\frac{\Sigma}{(2\pi)^{2}}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{\Sigma(4\varepsilon_{0})}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{2\cos(kr\cos\varphi)}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\frac{m_{h}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{h}}}-\frac{m_{e}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\right)\frac{k{\rm d}k{\rm d}\varphi}{m_{h}-m_{e}}$
@@
$=V_{\rm KR}(r,r_{\infty})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{2J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\left(\frac{m_{h}}{1+\sigma_{h,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}-\frac{m_{e}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}\right)\frac{\tilde{k}{\rm d}\tilde{k}}{\Delta m},$ (S137)
@@
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\left[\frac{m_{h}}{\Delta m}\frac{\pi(1+\sigma_{0})\tilde{\sigma}_{h,{\rm t}}\left(I_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)-L_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)\right)+2(\tilde{\sigma}_{h,{\rm t}}^{2}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)}{(1+\tilde{\sigma}_{h,{\rm t}}^{2})(\sigma_{0}^{2}+\tilde{\sigma}_{h,{\rm t}}^{2})}\right.$
@@
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})-E_{0}\left[\frac{m_{h}}{\Delta m}h_{\tilde{\sigma}_{h,{\rm t}}}(\varepsilon r/r_{\infty})-\frac{m_{e}}{\Delta m}h_{\tilde{\sigma}_{e,{\rm t}}}(\varepsilon r/r_{\infty})\right],$ (S151)
@@
$h_{\nu}(x)=\dfrac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\left[(1+\sigma_{0})\alpha\pi\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]+2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)\right].$ (S152)

### v2

@@ IV.1 Hamiltonian and unitary transformations
$V_{k}=ie\omega_{\rm t}\sqrt{\frac{\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}A\omega_{{\rm l},k}}}\frac{1}{\varepsilon+r_{\infty}k}.$ (S49)
@@ IV.2 Polaron energy minimization
$|V_{k}|^{2}$ $=\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{A(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)}}\frac{1}{(\varepsilon+r_{\infty}k)}.$ (S82)
@@
$\beta p$ $=\frac{A}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{A(4\varepsilon_{0})}\frac{1}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}+2\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)\right)\frac{\hbar k\cos\varphi k{\rm d}k{\rm d}\varphi}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{3}}$
@@
$=\frac{p^{2}}{2m_{e}}(1-\beta^{2})-\frac{A}{(2\pi)^{2}}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\left(1+\frac{\frac{\hbar}{m_{e}}pk\cos\varphi(1-\beta)}{\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}}+\frac{\frac{\hbar^{2}}{m_{e}^{2}}(1-\beta)^{2}(pk\cos\varphi)^{2}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)^{2}}\right)k{\rm d}k{\rm d}\varphi$
@@
$\chi_{e}=\frac{E_{0}}{\hbar\omega_{\rm t}}I_{m}(\sigma_{0},\sigma_{e,{\rm t}}),$ (S94)
@@
$\beta=\frac{\chi_{e}}{1+\chi_{e}},$ (S95)
@@
$E$ $=-E_{0}I_{\rm b}(\sigma_{0},\sigma_{e,{\rm t}})+\frac{p^{2}}{2m_{e}}\frac{1}{1+\chi_{e}}.$ (S96)
We can further proceed with approximate expressions for the integrals $I_{\rm b}$, $I_{m}$. One can note, that at the limit $\tilde{k}\gg 1$ one has $\sqrt{(1+\tilde{k})/(1+\sigma_{0}\tilde{k})}\approx 1/\sqrt{\sigma_{0}}$. We therefore perform the corresponding substitution, and get an expression which can be readily integrated, resulting in
Figure S1: The relative error of polaron binding energy integral and its analytic estimate $(I_{\rm b}-I_{\rm b}^{\rm approx})/I_{\rm b}$.
$I_{\rm b}^{\rm approx}=\frac{\sigma_{0}-1}{2(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})}\left(\frac{\pi(\sigma_{0}+1)\tilde{\sigma}_{e,{\rm t}}}{2\left(\tilde{\sigma}_{e,{\rm t}}^{2}+1\right)}+\frac{\left(\tilde{\sigma}_{e,{\rm t}}^{2}-\sigma_{0}\right)\log\tilde{\sigma_{e,{\rm t}}}}{\tilde{\sigma}_{e,{\rm t}}^{2}+1}-\frac{\sigma_{0}\log\sigma_{0}}{\sigma_{0}-1}\right),$ (S97)
where $\tilde{\sigma}_{e,{\rm t}}=\sigma_{e,{\rm t}}/\sigma_{0}^{1/4}$. Fig. S1 shows the relative error defined as $(I_{\rm b}-I_{\rm b}^{\rm approx})/I_{\rm b}$, demonstrating that the approximation works well at the limit $\sigma_{0}-1\ll 1$. Applying the same approximation, we get
$I_{m}^{\rm approx}=\frac{\sigma_{e,{\rm t}}^{2}(\sigma_{0}-1)}{8\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2})^{4}}\times$
$\left[2(\sigma_{0}^{6}+\tilde{\sigma}_{e,{\rm t}}^{6})-3\pi\sigma_{0}\tilde{\sigma}_{e,{\rm t}}\left(\sigma_{0}^{4}+\tilde{\sigma}_{e,{\rm t}}^{4}\right)-18\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}+\tilde{\sigma}_{e,{\rm t}}^{2}-\pi)+24\sigma_{0}^{2}\tilde{\sigma}_{e,{\rm t}}^{2}(\sigma_{0}^{2}-\tilde{\sigma}_{e,{\rm t}}^{2})\log\!\biggl{(}\sigma_{0}/\tilde{\sigma}_{e,{\rm t}}\biggr{)}\right].$ (S98)
IV.3 Landau-Pekar model
The polaron binding energy within the Landau-Pekar model can be found via the variational minimization of the following functional [63]:
$E_{e}^{\rm LP}=\left\langle\psi\left|\frac{\hat{p}_{e}^{2}}{2m_{e}}\right|\psi\right\rangle-\sum_{k}\frac{|V_{k}|^{2}|g_{k}|^{2}}{\hbar\omega_{{\rm l},k}},$ (S99)
Table 2: Binding energy of electron- and hole- polarons calculated calculated via different methods. Here $E_{e}^{\rm b}$ is the energy of Fröhlich electron-polaron, $E_{e}^{\rm LP}$ is the Landau-Pekar binding energy, calculated via Eq. (IV.3), and $E_{e}^{\rm LP*}$ is the Landau-Pekar binding energy calculated within the trivial limit of LO phonon dispersion in Ref. [43]. Similar results are shown for hole-polarons.
Material $E_{e}^{\rm b}$ (meV) $E_{e}^{\rm LP}$ (meV) $E_{e}^{\rm LP*}$ (meV) $E_{h}^{\rm b}$ (meV) $E_{h}^{\rm LP}$ (meV) $E_{h}^{\rm LP*}$ (meV)
h-BN 129.3 12.00 17.5 116.3 7.405 10.8
GaN 55.16 0.460 0.7 132.0 37.47 57.4
AlN 105.4 9.607 17.2 186.5 66.83 119.9
HfSe2 89.56 64.05 402.4 98.51 73.72 458.0
HfS2 123.2 82.04 469.0 155.1 115.8 646.0
ZrS2 142.9 80.02 349.0 133.7 71.34 313.1
where $g_{k}=\langle\psi\left|e^{i\mathbf{k}\cdot\mathbf{r}}\right|\psi\rangle$. The trial function is chosen in the 2D hydrogen-like form:
$\psi=\sqrt{\frac{2}{\pi}}\frac{1}{a}e^{-r/a},$ (S100)
where $a$ is a variational parameter. The kinetic energy is then readily evaluated as
$\left\langle\psi\left|\frac{\hat{p}_{e}^{2}}{2m_{e}}\right|\psi\right\rangle=\frac{\hbar^{2}}{2m_{e}a^{2}},$ (S101)
and
$g_{k}=\frac{1}{\left(1+(ak)^{2}/4\right)^{3/2}}.$ (S102)
The explicit expression for the effective potential is
$\frac{|V_{k}|^{2}}{\hbar\omega_{{\rm l},k}}=\frac{1}{\hbar\omega_{{\rm l},k}}\frac{e^{2}\omega_{\rm t}^{2}\hbar(r_{0}-r_{\infty})}{4\varepsilon_{0}A\omega_{{\rm l},k}}\frac{1}{(\varepsilon+r_{\infty}k)^{2}}=\frac{e^{2}(r_{0}-r_{\infty})}{4\varepsilon_{0}A}\frac{1}{(\varepsilon+r_{\infty}k)(\varepsilon+r_{0}k)}.$ (S103)
Then one has
$E_{e}^{\rm LP}$ $=\frac{\hbar^{2}}{2m_{e}a^{2}}-\frac{A}{(2\pi)^{2}}\int\limits_{0}^{\infty}\frac{e^{2}(r_{0}-r_{\infty})}{4\varepsilon_{0}A}\frac{1}{(\varepsilon+r_{\infty}k)}\frac{1}{(\varepsilon+r_{0}k)}\frac{2\pi k{\rm d}k}{\left(1+(ak)^{2}/4\right)^{3}}$
@@ IV.3 Landau-Pekar model
$=\frac{\hbar^{2}}{2m_{e}a^{2}}-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{1}{(1+\tilde{k})}\frac{1}{(1+\sigma_{0}\tilde{k})}\frac{\tilde{k}{\rm d}\tilde{k}}{\left(1+\left(\tilde{k}\frac{\varepsilon a}{2r_{\infty}}\right)^{2}\right)^{3}}.$ (S104)
Variational minimization of the last expression gives the polaron binding energy in the Landau-Pekar treatment. The results are shown in Table 2. One can find essentially smaller binding energy obtained within Landau-Pekar approach, in line with the 3D case [54]. This difference is especially dramatic for weakly polar materials, where $r_{0}-r_{\infty}<r_{\infty}$, such as h-BN, GaN, AlN. We additionally address the results of Ref. [43], where the effective potential for Landau-Pekar model is
$\frac{V_{k}^{2}}{\hbar\omega}=\frac{e^{2}(\epsilon_{0}-\epsilon_{\infty})dq_{c}^{2}}{16\pi\varepsilon_{0}}\frac{2\pi}{A}\frac{1}{(k+q_{c})^{2}},$ (S105)
with $q_{c}=1/[d\epsilon_{\infty}/2-d/(4\epsilon_{\infty})]$. Substituting $d\epsilon_{\infty}/2=r_{\infty}$, $d\epsilon_{0}/2=r_{0}$, we get $q_{c}=1/[r_{\infty}-r_{\infty}/(2\epsilon_{\infty}^{2})]$, and
$\frac{V_{k}^{2}}{\hbar\omega}=\frac{e^{2}(r_{0}-r_{\infty})}{4\varepsilon_{0}A}\frac{1}{[1+kr_{\infty}(1-1/(2\epsilon_{\infty}^{2}))]^{2}}.$ (S106)
Now, taking into account the dielectric constant of surrounding media $\varepsilon=1$, we get
$E_{e}^{\rm LP*}$ $=\frac{\hbar^{2}}{2m_{e}a^{2}}-\frac{A}{(2\pi)^{2}}\int\limits_{0}^{\infty}\frac{e^{2}(r_{0}-r_{\infty})}{4\varepsilon_{0}A}\frac{1}{[1+kr_{\infty}(1-1/(2\epsilon_{\infty}^{2}))]^{2}}\frac{2\pi k{\rm d}k}{\left(1+(ak)^{2}/4\right)^{3}}$
$=\frac{\hbar^{2}}{2m_{e}a^{2}}-\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{1}{\left[1+\tilde{k}(1-1/(2\epsilon_{\infty}^{2}))\right]^{2}}\frac{\tilde{k}{\rm d}\tilde{k}}{\left(1+\left(\tilde{k}\frac{a}{2r_{\infty}}\right)^{2}\right)^{3}}.$ (S107)
The variational minimization of Eq. (IV.3) allows to exactly reproduce polaron binding energies reported in Ref. [43]. One can thus see that neglecting the dispersion of 2D LO phonons has a dramatic impact on polaron binding energy, leading to overestimation for strongly polar materials with $r_{0}\gg r_{\infty}$.
@@ V Exciton states
$A_{\mathbf{k}}=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{e}}\right)},\quad B_{\mathbf{k}}=\frac{V_{k}^{*}}{\left(\hbar\omega_{{\rm l},k}+\frac{\hbar^{2}k^{2}}{2m_{h}}\right)}.$ (S142)
@@
$\frac{A}{(2\pi)^{2}}\frac{e^{2}\omega_{\rm t}\hbar(r_{0}-r_{\infty})}{A(4\varepsilon_{0})}\int\limits_{0}^{\infty}\int\limits_{0}^{2\pi}\frac{2\cos(kr\cos\varphi)}{\sqrt{(\varepsilon+r_{0}k)(\varepsilon+r_{\infty}k)^{3}}}\left(\frac{m_{h}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{h}}}-\frac{m_{e}}{\hbar\omega_{\rm t}\sqrt{\frac{\varepsilon+r_{0}k}{\varepsilon+r_{\infty}k}}+\frac{\hbar^{2}k^{2}}{2m_{e}}}\right)\frac{k{\rm d}k{\rm d}\varphi}{m_{h}-m_{e}}$
@@
$=\boxed{V_{\rm KR}(r,r_{\infty})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\int\limits_{0}^{\infty}\frac{2J_{0}(\tilde{k}\varepsilon r/r_{\infty})}{(1+\sigma_{0}\tilde{k})(1+\tilde{k})}\left(\frac{m_{h}}{1+\sigma_{h,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}-\frac{m_{e}}{1+\sigma_{e,{\rm t}}^{2}\tilde{k}^{2}\sqrt{\frac{1+\tilde{k}}{1+\sigma_{0}\tilde{k}}}}\right)\frac{\tilde{k}{\rm d}\tilde{k}}{\Delta m}},$ (S145)
@@
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})+\frac{e^{2}}{4\pi\varepsilon_{0}r_{\infty}}\frac{\sigma_{0}-1}{2}\left[\frac{m_{h}}{\Delta m}\frac{\pi(1+\sigma_{0})\tilde{\sigma}_{h,{\rm t}}\left(I_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)-L_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)\right)+2(\tilde{\sigma}_{h,{\rm t}}^{2}-\sigma_{0})K_{0}\left(\frac{\varepsilon r}{\tilde{\sigma}_{h,{\rm t}}r_{\infty}}\right)}{(1+\tilde{\sigma}_{h,{\rm t}}^{2})(\sigma_{0}^{2}+\tilde{\sigma}_{h,{\rm t}}^{2})}\right.$
@@
$V_{\rm eff}(r)$ $\approx V_{\rm KR}(r,r_{0})+E_{0}\left[\frac{m_{h}}{\Delta m}\Phi_{\tilde{\sigma}_{h,{\rm t}}}(\varepsilon r/r_{\infty})-\frac{m_{e}}{\Delta m}\Phi_{\tilde{\sigma}_{e,{\rm t}}}(\varepsilon r/r_{\infty})\right],$ (S159)
@@
$\Phi_{\nu}(x)=\dfrac{\sigma_{0}-1}{2(1+\nu^{2})(\sigma_{0}^{2}+\nu^{2})}\left[(1+\sigma_{0})\nu\pi\left[I_{0}\left(x/\nu\right)-L_{0}\left(x/\nu\right)\right]+2(\nu^{2}-\sigma_{0})K_{0}\left(x/\nu\right)\right].$ (S160)